RC Meeting and Run Plan, Sunday, February 17

From HallCWiki
Jump to: navigation, search

Main | Online | RC | Run Plan | Official Website | Redmine | Index | Nav
Jpsi007.png

RC Meeting Information

RC Daily Meetings are at 4:30 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Current Status

  • We are currently in Restricted Access.
  • Beam was running on J/Psi production #1-1 LH2 setting at 45uA.
  • We had a sudden loss of LCW to Hall C which shut down all HMS magnets and SHMS HB. The water hose in the new HMS Q2 power supply popped off and subsequently generated a waterfall from the top of the HMS carriage to the floor. We shut off the water to the HMS Q2, but the LCW system for all the halls is currently being purged and filled and is expected to return in a few hours (noon-ish). In the meantime, floor has been mostly dried out by the HMS, fans are on the HMS Q1 power supply which had some water on its boards. The magnets appear to be fine, but we need to check out the HMS Q1 power supply. Once LCW is restored, Steve and Jack will check out the magnet power supplies to determine if we can run with them. The plan will be to proceed with J/Psi run plan. Preliminary estimate is that we will be locking up at or by the end of the Day Shift.
  • Here's pictures and videos that capture the water: https://logbooks.jlab.org/entry/3654945

Current Run Plan

  • More x>1 running
  • Detailed Instructions:
  • Alternate:
    • 1 hour on 10B4C, ps2=0 (keep rates below 5kHz)
    • 1 hour on 11B4C, ps2=0 (keep rates below 5kHz)
    • 1.5 hours on LD2, ps2=2 (keep rates below 5kHz), 30 uA

After 2 cycles - add carbon to the list.

  • Take 1 dummy run, 1 hour (keep rates on ps2 below 5kHz)
  • End runs every 30 minutes.
  • Keep track of yields per target on each 50k replay using Eric's macro
    • In /home/cdaq/hallc-online/hallc_replay/macros/ execute yield_xgt1.C
root -l
.x yield_xgt1.C (runNum, 50000)

Notices in Effect for Shift Crew

  • If there is a magnet Interlock, check the power supplies on the camera for smoke and floor for water before re-setting the interlock!!!
  • HMS Q2 readback current box does not work. The current listed above it is the most useful indication that it is on.
  • DO NOT RAMP HMS Q2 to 0! If it reaches 10A, just turn it off. Let the magnet settle for about 5 minutes at the set current before changing it.
  • We have no control over polarity for HMS magnets now. They don't have meaning in the GUI, either. It is currently set for positron/proton running (Q1-, Q2+ on hall probes)
  • If rotation cannot be Reset, call the RC, then call Ops for an escorted/controlled access.

Current Issues

  • SHMS vacuum is at 3E-7, but the indicator in the magnet GUI is showing an error with E-11. Probably a read back error, but the gauge appears to be working.

Planned Accesses to the Hall

Planned Accelerator Configuration Changes

Special Requests